HiTech Products manufacures three types of DVD players. Economy, Standard and Deluxe. The company, which uses activity-based costing, has identified five activities (and relative cost drivers). Each activity, its budgeted cost and related cost driver is identified below:

Activity Cost Cost Driver
Material Handling $225,000 Number of parts
Material insertion 2,475,000 Number of parts
Automated machinery 840,000 Machine hours
Finishing 170,000 Direct lbor hours
Packaging 170,000 Order shipped
Total 3,880,000

The following information pertains to each product line of DVD players for next year:

Economy Standard Deluxe
Units to be produced 10,000 5,000 2,000
Orders to be shipped 1,000 500 200
Number of parts per unit 10 15 25
Machine hours per unit 1 3 5
Labor hours per unit 2 2 2

Assuming that the company is working at capacity:

1. What is the cost application rate for the material handing activity?

2. What is the cost application rate for the packaging activity?

3. Under an activity based costing system, what is the per unit cost of Economy? And for Deluxe?

4. Assume that HiTech is using a volume-based costing system, and the preceding manufacturing costs are applied to all products based on direct labor hours. How much of the preceding cost is assigned to Deluxe?
Answer #1

1.

Material handling cost = 225,000

No. of parts

Economy = 10000 x 10 = 100,000

Standard = 5000 x 15 = 75000

Deluxe = 2000 x 25 = 50000

Total no. of parts = 100,000 + 75,000 + 50000 = 225,000

Cost Application rate for the material handling activity

= Material handling cost / Total No. of parts * No. of parts per activity

Economy = 225,000 / 225,000 * 100,000 = 100,000

Standard = 225,000 / 225,000 * 75,000 = 75,000

Deluxe = 225,000 / 225,000 * 50,000 = 50,000

2. Cost application rate for packaging activity

Packaging cost / No. of orders shipped

Economy = 170,000 / 1,700 * 1000 = 100,000

Standard = 170,000 / 1,700 * 500 = 50,000

Deluxe = 170,000 / 1,700 * 200 = 20,000

3. Calculation of per unit cost

Economy Deluxe

Material Handling 100,000    50,000

Material insertion 1,100,000 5,50,000

Automated machinery 240,000      240,000

Finishing 100,000 20,000

Packaging     100,000 20,000

--------------------- ---------------------------

Total Cost (A) 1,640,000 880,000

Units produced (B) 10,000 2,000

Per unit cost (A) / (B) 164 per unit 440 per unit